OpenStudy (anonymous):

what percent of 300 is 40?

OpenStudy (anonymous):

x/100 = 40/300

OpenStudy (disco619):

\[\frac{ 40 }{ 300 } \times 100 = x\]

OpenStudy (disco619):

First solve the Fraction then multiply by Hundred. Solved!

OpenStudy (the_fizicx99):

Make a proportion. x / 100 = 40/300 Cross mulitply

OpenStudy (the_fizicx99):

Or really just solve 40/300 and what ever that is just multiply it by 100, per hundred to get your percent